OpenStudy (anonymous):

Prove: sin θ - sin θ cos2 θ = sin3 θ. I have to show all work

OpenStudy (anonymous):

@ganeshie8 @kc_kennylau @ranga @shamil98 @AriPotta @ganeshie8 @TuringTest @thomaster

OpenStudy (anonymous):

@Tazmann @ganeshie8 @TuringTest @thomaster @zimmah

OpenStudy (turingtest):

is that \[ \sin \theta - \sin \theta \cos(2\theta) = \sin(3 \theta)\]or \[ \sin \theta - \sin \theta \cos^2\theta = \sin^3 \theta\]or some combination thereof?

OpenStudy (anonymous):

I will attach a pic of the question

OpenStudy (turingtest):

why couldn't you have just said "the second one you wrote" ?

OpenStudy (anonymous):

cos I wasnt sure it was the second one you wrote

OpenStudy (anonymous):

just to avoid error

OpenStudy (turingtest):

try factoring the left side

OpenStudy (anonymous):

how do i do that?

OpenStudy (turingtest):

could you factor x-xy ?

OpenStudy (anonymous):

x (1 - y)

OpenStudy (anonymous):

?

OpenStudy (turingtest):

yes, now let sin(t)=x and cos^2(t)=y and factor it the same way

OpenStudy (raden):

after you factor out the left side, then use the identity : sin^2 θ + cos^2 θ = 1

OpenStudy (anonymous):

I cant figure it out.

OpenStudy (anonymous):

What should me equation look like?

OpenStudy (anonymous):

*my

OpenStudy (turingtest):

x=sin(t) y=cos^2(t) substitute to get x-xy factor this, as you showed you are capable of, and substitute back in for x and y

OpenStudy (anonymous):

So sin(t) - sin(t)cos^2(t)?

OpenStudy (turingtest):

that is still in the form x-xy factor this and what do you get?

OpenStudy (anonymous):

how do i go about factoring this?

OpenStudy (turingtest):

you already did it once above

OpenStudy (turingtest):

I said "can you factor x-xy ?" and you wrote the answer reread above

OpenStudy (anonymous):

so sin(t) (1 - cos(t^2))

OpenStudy (turingtest):

almost, but\[\cos^2\theta=(\cos\theta)^2\ne\cos(\theta^2)\]

OpenStudy (turingtest):

you have confused the two

OpenStudy (anonymous):

I really don't get it. can you factor it for me?

OpenStudy (turingtest):

you factored it fine, but there never was a cos(t^2) in your problem

OpenStudy (turingtest):

there is a difference between cos^2(t) and cos(t^2), you need to understand that

OpenStudy (anonymous):

yeah but i don't know how to factor it when cos^2(t)

OpenStudy (turingtest):

I just told you twice how to factor it, and you did factor it, and then, for some reason I don't understand, you changed cos^2(t) to cos(t^2) aside from that it's fine so far

OpenStudy (anonymous):

You can't just keep on telling me how to factor it if i don't know how! x-xy is one thing, that problem is quite another

OpenStudy (turingtest):

actually they are identical x=sin(t) y=cos^2(t) now we can forget what x and y stand for and substitute, giving x-xy=x(1-y) now just make the *exact* same substitution in reverse to go back to sin and cos